Let f(x)= -x^2+2x-3 and g(x)=-x^2+4x Find each of the following.
a. f(-2)
b. g(3)
c. g(-k)
d. g(3m)
e. g(k-5)
f. f(3-p)
I was able to get the following.....
a. 5(-2)-3
-10-3
=-13
b. -3^2+4(3)
-3^2+12
-9+12
=3
c. -k^2+4(-k)
=-k2+4k
d. -3m^2+4(3m)
-3^2+12m
=-9m2+12m
THe following is all that I could get for letter E but I dont think it is right can you please help? The answer is supposed to be -k^2+14k-45 but I cant get there
e. (-k-5)^2=4(k-5)
(k+5)^2+4(k-5)
(k+5)^2=4k+20

The following is all I could get for letter F. The answer is supposed to be 12-5p.
f. 5(3-p)-3
15-5p-3
10p-3
thank you for your assistance :)

Answer by stanbon(48535)

Let f(x)= -x^2+2x-3 and g(x)=-x^2+4x Find each of the following.
a. f(-2) = -(-2)^2+2(-2)-3 = -4-4-3=-11
b. g(3) = -(3)^2+4*3 = -9+12 = 3
c. g(-k) = -(-k)^2+4*(-k) = -k^2-4k
d. g(3m) = -(3m)^2+4(3m) = -9m^2+12m
e. g(k-5)= -(k-5)^2+4(k-5) = -k^2+10k-25+4k-20 = -k^2+14k=45
f. f(3-p)= -(3-p)^2+2(3-p)-3 = -9+6p-p^2+6-2p-3= -p^2+4p-12
==============
Cheers,
Stan H.